explain clearly the concept of weighted average cost of capital (wacc), how it can be computed step by step, and how it can be used?
Answer #1

Ans) Cost of equity is minimum annual  rate of earning on equity funds expected by the owners of business.In absolute terms this is the minimum profit after interest and tax a business must earn.For example if value of equity of a business is $100 and its cost of equity is 20 % than the buainess must earn 20% of 100 or $20 after interest and tax.In other words coat of equity is the cost paid to obtain equity as a source of finance .

Cost of debt is the cost associated for using debt as a source of finance or is the cost of servicing a debt.The cost of debt service consists of coupons , redemption premium and other payments for debt service made by the borrower.

Weighted average cost of capital is the weighted average of cost of debt and cost of equity .Here weight used are proportions of equity and debt in market value.

We take an example to better understand the Weighted average cost of capital(WACC).

Let suppose cost of equity is 20% and cost of debt is 15%.Tax rate is 40% if the value of equity is $40000 the business should earn $ 8000 ( 20% of $ 40000) after interest and tax to satisfy the equity holders.If the value of debt is $60000 the company should pay $ 9000 ( 15% of 60000) as interest.Hence interest after tax $ 5400 ( 60% of 9000).

The weighted average cost of capital is calculate as follows :8000+5400 = 13400 and as a rate it is =( 13400/100000 )*100 ; ( value of equity + value of debt ,which is 40000+60000 = 100000)

=13.4% .The business cannot afford to finance any project with this fund , unless the project earns 13.4% or more all projects earning leas rhan 13.4 % must be rejected hence it is very useful to dind out the cut-off rate that is the rate of return that cuts off the unacceptable projects.